OpenStudy (anonymous):

Determine whether there exists a real number x satisfying sqrt(x-2) = 3 - 2*sqrt(x). If there is, determine this number. So I square both sides which gives x-2 = 9 - 12*sqrt(x) + 4x => 12*sqrt(x) = 11 + 3x Squaring both sides gives x = (121+66x+9x^2)/12 which clearly doesn't have any real solutions, as the book's answer gives. However, I entered the original equation into WolframAlpha and it returned a real solution: 13/3 - 4/sqrt(3) Who's wrong?

OpenStudy (anonymous):

sqrt(x-2) = 3 - 2*sqrt(x). square both sides which gives x-2 =9 - 12*sqrt(x) + 4x 12*sqrt(x) = 11 + 3x square both side gives x = (121+66x+9x^2)/144

OpenStudy (anonymous):

got your mistake?

OpenStudy (anonymous):

Ah! I just caught my error! Yes thank you I just realized I didn't square the denominator in the last bit.

OpenStudy (anonymous):

Welcome